OpenStudy (anonymous):

in triangle DEF DE=5 and m

OpenStudy (anonymous):



OpenStudy (mathstudent55):



OpenStudy (mathstudent55):

For angle D of 55 degrees, leg FE is opp or adj? For angle D, leg DE is opp or adj?

jhonyy9 (jhonyy9):

angle D or angle E is right angle ?

OpenStudy (anonymous):

i ot 47

jhonyy9 (jhonyy9):

sorry E not is on the right place

jhonyy9 (jhonyy9):

tan D = FE / 5

jhonyy9 (jhonyy9):

tan 55 = 1,55 1,55=FE/5 FE = 1,55*5 = 7,75 FE=7,75

jhonyy9 (jhonyy9):

you have any choises ?

OpenStudy (anonymous):

its 47

jhonyy9 (jhonyy9):

so this mean that there is somthing missed or confussed with D or E or i dont know

jhonyy9 (jhonyy9):

so this 47 is what

jhonyy9 (jhonyy9):

47 wann being the length of side FE ???

jhonyy9 (jhonyy9):

@Luis_Rivera do you understand it ?

OpenStudy (anonymous):

to give an appropiate answer, the question must be stated appropiatly. please do a better job.